OpenStudy (anonymous):

If k is an integer, then k^2+3k is even. Proof by cases

OpenStudy (anonymous):

or is there only two cases, when k is odd or k is even

terenzreignz (terenzreignz):

Well, since we're dealing with parity anyway, your instincts should tell you to consider the case where k is odd, or k is even.

terenzreignz (terenzreignz):

Already got it?

OpenStudy (anonymous):

Can you help me out with another?

terenzreignz (terenzreignz):

Will try :)

OpenStudy (anonymous):

If a and b are aribitary real numbers, then a*b<= (a^2 + b^2 )/2

OpenStudy (anonymous):

I was thinking of starting by saying 0<=a^2+ b^2-2a*b, but i dont know if thats gonna get me anywhwere.

terenzreignz (terenzreignz):

hmm a² - 2ab + b² = (a-b)²

OpenStudy (anonymous):

hmm....

OpenStudy (anonymous):

its not coming to me, i dont know what to do with 0<=(a-b)^2

terenzreignz (terenzreignz):

You do agree that (a-b)² is always nonnegative, right? as it's a square? :)

OpenStudy (anonymous):

oh right, i didnt even think about....didnt get enough sleep last night i guess

terenzreignz (terenzreignz):

so, if a and b are real numbers, it's always true that 0 <= (a - b)² Expanding... 0 <=a² - 2ab + b² And the rest is history :)

OpenStudy (anonymous):

okay, but then from here: 0 <=a² - 2ab + b² im still lost a little bit, can you explain the concept, from here: 0 <=a² - 2ab + b², how to we aregure that a*b<= (a^2 + b^2 )/2

terenzreignz (terenzreignz):

Well, using the addition property of inequality, we can add 2ab to both sides, giving us 2ab <= a² + b² Now using the multiplication property of inequality, we can multiply both sides by 1/2, and since it's a positive number, the inequality doesn't change... ab <= (a² + b²)/2 voila

OpenStudy (anonymous):

i think i finally got it, thanks :)

terenzreignz (terenzreignz):

No problem :)

OpenStudy (anonymous):

Let x be an arbritary real number. If |x-2|>4, then |x|>2. Prove.

OpenStudy (anonymous):